OpenStudy (anonymous):

Complete the square on the quadratic y=-2x^2+3x+6

OpenStudy (anonymous):

please show your work

OpenStudy (anonymous):

I have gotten this far y=-2(x^2+3x)+6 Y=-2(x^2+? what at this point

hero (hero):

y = -2(x^2+[3x/-2]-3)

hero (hero):

When you factor out a number from each term, you divide that number by each term.

OpenStudy (anonymous):

So is this as far as I go then i solve for th y intercept?

hero (hero):

I think you were doing it right. You just needed to understand how to find the middle term.

OpenStudy (anonymous):

I just don't know what to do after I find that term. I seem to be having a mental block

hero (hero):

You just have to find the complete square by adding (b/2)^2 and then reducing the terms to the form (x-a)^2